OpenStudy (anonymous):

a solid sphere of radius=r is melted &rec into d shape of a solid cone of height 4r.d r of base of cone

OpenStudy (callisto):

volume of the sphere = 4/3Π(r^3), vol of cone = vol. of sphere, 1/3Π(r^2)h = 4/3Π(r^3), 1/3Π(r^2)*(4r) = 4/3Π(r^3), r of the cone = r of the sphere

OpenStudy (anonymous):

h=r

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thanks